OpenStudy (anonymous):

8k^2-19k+9 factor if possible and show work

myininaya (myininaya):

Find factors of 8(9) that add up to be -19 if there aren't any then say not factorable over the integers

OpenStudy (anonymous):

yeah it shld b -18k

OpenStudy (anonymous):

how are you getting -18k??

myininaya (myininaya):

-8(-9)=72 but -8-9=17 so that will not work -4(-18)=72 but -4-18=-22 -12(-6)=72 but -12-6=-18 -24(-3)=72 but -24-3=-27 if you can't find any factors of a*c that add up to be b, then say not factorable but i have to go

OpenStudy (anonymous):

i dont get that

OpenStudy (anonymous):

im sayin d question myt b wrng....its only factorable into irrational nos

OpenStudy (anonymous):

so then it would be a prime trinomial??

OpenStudy (anonymous):

cause it cant be factored by integers

OpenStudy (anonymous):

wts a prime trinomial?

OpenStudy (anonymous):

book says a plynomial that cant be written as a product of two plynomials with integeral coefficients